OpenStudy (anonymous):

Use the first and second derivative tests to find the local min and max of the function. :D

OpenStudy (anonymous):

\[1+3x^2-2x^3\]

OpenStudy (anonymous):

\[f'(x)=-6x^2+6x\]

OpenStudy (anonymous):

why both? one should be enough....

OpenStudy (anonymous):

\[6x(-x^2+1)=0\]

OpenStudy (anonymous):

im not sure how to do the second derivative test...

OpenStudy (anonymous):

critical points are 0, 1

OpenStudy (anonymous):

f(0)=min and f(1)=max

OpenStudy (anonymous):

\[f''x=-12x+6\]

OpenStudy (anonymous):

\(\large f'(x)=6x- 6x^2=6x(1-x)\)

OpenStudy (anonymous):

x=1/2 for the second one...

OpenStudy (anonymous):

oh ya thats right,

OpenStudy (anonymous):

evaluate the cp on the second derivative... what do you get?

OpenStudy (anonymous):

x=1/2

OpenStudy (anonymous):

i mean critical numbers, not critical points.... evaluate the second derivative on those cr. #s... what do you get?

OpenStudy (anonymous):

f''(0) = ??? f''(1) = ???

OpenStudy (anonymous):

so i take 0 and 1 and plug them in?

zepdrix (zepdrix):

Calle, I'm not sure how you figured out that the 0 was your min, I guess you used your smarty brain. :D Another way to determine which one is a min, and which is a max, (the way that they want you to try), is by doing the second derivative test. Plug your critical points into the second derivative, If the result is positive, then that part of the graph is concave up, meaning you have a minimum point. If the result is negative, then that part of the graph is concave down, meaning you have a maximum point. Yah try that :D

OpenStudy (anonymous):

can't i find the max and min from plugging in the first derivative critical points into the origional equation?

zepdrix (zepdrix):

Ummmmmmmmmmm I think that will USUALLY work, yes. But they didn't want you to determine it that way, they want you to practice using the second derivative test! :)

OpenStudy (anonymous):

so, the second derivative is -12x+6

zepdrix (zepdrix):

Because what comes next in calc is, being able to graph a function like that without a calculator. So you really want to know where it curves upward, and downward. So you need inflection points. Using the second derivative test, i think, is just to prepare you for the next step :D

zepdrix (zepdrix):

So determine if f''(0) is + or - and also check f''(1) :D

OpenStudy (anonymous):

how can i find the max and min from the second derivative?

zepdrix (zepdrix):

DO what i just told you to do ! >:O lol

zepdrix (zepdrix):

f''(0) = -12(0) + 6 Is that positive or negative? :D

OpenStudy (anonymous):

f(0) is the max f(1) is the min

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

the first and second derivative contradict each other? :(

zepdrix (zepdrix):

 f(0)=1+0-0=1 f(1)=1+3-2=2 I think that agrees with what you get when you plug it into the function :O Did you punch some numbers wrong maybe? <:O Maybe I did :D I'll take another look.